OpenStudy (anonymous):

What are the solutions to 6x^2=18x? A)x=6 x=3 B)x=6 x=-3 c)x=0 x=-3 D)x=0 x=3

OpenStudy (anonymous):

just plug them in

OpenStudy (anonymous):

Is there another way?

OpenStudy (mathmale):

There surely is another way. Solve this quadratic equation: 6x^2=18x Re-write it as 6x^2-18x=0. Factor this equation. Set the factors equal to zero, separately. What are the resulting x values? Check your solutions by substituting them back into the original equation.

OpenStudy (anonymous):

yes, subtract 18x so that the equation is equal to 0 (6x^2-18x=0) then factor out 6x to get x-3=0 then add three and you get 3. then take that 6x you factored with and set that equal to zero and get zero. so the answer is (0,3)

OpenStudy (anonymous):